# ResSVD: Residual Compensated SVD for Large Language Model Compression

### v1

@@ 3 ResSVD
The framework of ResSVD is illustrated in Figure 1. We first perform SVD to compress the last few layers of the model, while ensuring that the overall compression ratio is satisfied, and compute the corresponding residual matrix. A second SVD is then applied to the residual matrix to obtain its low-rank approximation. The two truncated matrices are subsequently combined to construct the final compressed weight matrices. In Section 3.1, we first provide a theoretical analysis demonstrating how residual compensation effectively reduces truncation loss. Subsequently, in Section 3.2, we explain the benefits of partial-layer compression in mitigating error propagation. The pseudocode of ResSVD is shown in Algorithm 1, and the pseudocodes for residual compensation and partial-layer compression are provided in Appendix.
@@ 4.2 Results
We evaluate the overall performance of ResSVD from three aspects: ① Effectiveness under different compression ratios (ranging from 20% to 60% in increments of 10%), ② Scalability when applied to larger-scale models, and ③ Generalizability across diverse LLM families. In addition, qualitative results, such as generated contents produced by compressed models, are provided in Appendix, offering a more intuitive and direct comparison.
@@
Performance under different compression ratios. We evaluate the performance of LLaMA-7B [36] and LLaMA-2-7B [37] compressed by ResSVD, conventional SVD, and existing post-training baselines under compression ratios ranging from 20% to 60% across 10 benchmark datasets. The results for LLaMA-7B are presented in Table 1, while those for LLaMA-2-7B are provided in Appendix. On the three language modeling datasets, WikiText-2 [27], PTB [26], and C4 [30], ResSVD consistently outperforms all baselines across evaluated compression ratios. In particular, on PTB and C4, the improvements are more pronounced, suggesting that ResSVD exhibits stronger generalization capability. More importantly, under a relatively high compression ratio (e.g., 50%), ResSVD still achieves substantial perplexity reductions compared to the existing best-performing baseline SVD-LLM [39]: 7% on WikiText-2, 14% on PTB, and 41% on C4. This demonstrates that ResSVD maintains an obvious performance advantage even under aggressive compression. In addition, on seven common sense reasoning datasets, ResSVD surpasses the existing best baseline on the majority of datasets, with an average accuracy gain of up to 9%, and a minimum improvement of 3% improvement, further highlighting its robustness and overall effectiveness.
@@ 4.3 Ablation Study
Effectiveness of residual compensation and partial-layer compression. As shown in Table 4, we conduct ablation studies to evaluate the individual contributions of residual compensation (REC) and partial-layer compression (PLC). Specifically, the left table presents the results for REC under multiple compression ratios. Incorporating REC consistently reduces perplexity across all settings, and ResSVD achieves the best performance across all datasets and baselines. The right table reports the results for PLC. Without PLC, perplexity on WikiText-2 [27] increases slightly, indicating the effect of cumulative error. In contrast, applying PLC consistently enhances performance, further amplifying the benefits of REC. These results confirm that PLC is essential for mitigating error propagation. More ablation results are provided in Appendix.
Impact of layer selection strategy. We compare the performance of LLaMA-7B [36] under 20% compression ratio with different layer selection strategies. The results are shown in Table 4.4. We observe that the performance of the compressed model varies significantly depending on which layers are selected for compression. The worst performance is observed when compressing the first 8 layers, primarily because the compression ratio per layer becomes excessively high, and error accumulation begins from the first layer. Across all configurations, compressing the last 16 layers yields the best overall performance. However, as more layers are compressed, the performance gradually declines, which can be attributed to increased cumulative approximation error.
@@ 4.4 Efficiency Results
Table 4: Ablation results of residual compensation (REC) and partial-layer compression (PLC) on LLaMA-7B [36] under compression ratios 20% and 30%. More ablation results are in Appendix.

### v2

@@ 3 ResSVD
The framework of ResSVD is illustrated in Figure 1. We first perform SVD to compress the last few layers of the model, while ensuring that the overall compression ratio is satisfied, and compute the corresponding residual matrix. A second SVD is then applied to the residual matrix to obtain its low-rank approximation. The two truncated matrices are subsequently combined to construct the final compressed weight matrices. In Section 3.1, we first provide a theoretical analysis demonstrating how residual compensation effectively reduces truncation loss. Subsequently, in Section 3.2, we explain the benefits of partial-layer compression in mitigating error propagation. The pseudocode of ResSVD is shown in Algorithm 1, and the pseudocodes for residual compensation and partial-layer compression are provided in Appendix A.1.
@@ 4.2 Results
We evaluate the overall performance of ResSVD from three aspects: ① Effectiveness under different compression ratios (ranging from 20% to 60% in increments of 10%), ② Scalability when applied to larger-scale models, and ③ Generalizability across diverse LLM families. In addition, qualitative results, such as generated contents produced by compressed models, are provided in Appendix A.4, offering a more intuitive and direct comparison.
@@
Performance under different compression ratios. We evaluate the performance of LLaMA-7B [38] and LLaMA-2-7B [39] compressed by ResSVD, conventional SVD, and existing post-training baselines under compression ratios ranging from 20% to 60% across 10 benchmark datasets. The results for LLaMA-7B are presented in Table 1, while those for LLaMA-2-7B are provided in Appendix A.2. On the three language modeling datasets, WikiText-2 [29], PTB [28], and C4 [32], ResSVD consistently outperforms all baselines across evaluated compression ratios. In particular, on PTB and C4, the improvements are more pronounced, suggesting that ResSVD exhibits stronger generalization capability. More importantly, under a relatively high compression ratio (e.g., 50%), ResSVD still achieves substantial perplexity reductions compared to the existing best-performing baseline SVD-LLM [41]: 7% on WikiText-2, 14% on PTB, and 41% on C4. This demonstrates that ResSVD maintains an obvious performance advantage even under aggressive compression. In addition, on seven common sense reasoning datasets, ResSVD surpasses the existing best baseline on the majority of datasets, with an average accuracy gain of up to 9%, and a minimum improvement of 3% improvement, further highlighting its robustness and overall effectiveness.
@@ 4.3 Ablation Study
Effectiveness of residual compensation and partial-layer compression. As shown in Table 4, we conduct ablation studies to evaluate the individual contributions of residual compensation (REC) and partial-layer compression (PLC). Specifically, the left table presents the results for REC under multiple compression ratios. Incorporating REC consistently reduces perplexity across all settings, and ResSVD achieves the best performance across all datasets and baselines. The right table reports the results for PLC. Without PLC, perplexity on WikiText-2 [29] increases slightly, indicating the effect of cumulative error. In contrast, applying PLC consistently enhances performance, further amplifying the benefits of REC. These results confirm that PLC is essential for mitigating error propagation. More ablation results are provided in Appendix A.3.
Impact of layer selection strategy. We compare the performance of LLaMA-7B [38] under 20% compression ratio with different layer selection strategies. The results are shown in Table 5. We observe that the performance of the compressed model varies significantly depending on which layers are selected for compression. The worst performance is observed when compressing the first 8 layers, primarily because the compression ratio per layer becomes excessively high, and error accumulation begins from the first layer. Across all configurations, compressing the last 16 layers yields the best overall performance. However, as more layers are compressed, the performance gradually declines, which can be attributed to increased cumulative approximation error.
@@ 4.4 Efficiency Results
Table 4: Ablation results of residual compensation (REC) and partial-layer compression (PLC) on LLaMA-7B [38] under compression ratios 20% and 30%. More ablation results are in Table 8.
@@ 5 Conclusion
Appendix A Appendix
A.1 Pseudocode
Algorithm 2 and Algorithm 3 present the pseudocode for partial-layer compression and residual compensation, respectively. Algorithm 2 identifies the optimal number of last layers $k^{\star}$ to compress, along with their corresponding layer compression ratio $R_{l}^{\star}$. Specifically, for a model with $N$ layers, we iterate over candidate values of $k$ using a step size $s$ and compute the associated $R_{l}$ for each. For each candidate configuration, we invoke Algorithm 3 to perform the compression, after which we compute the final-layer error relative to the original model. The configuration yielding the lowest final-layer error is selected as the optimal compression setting.
Algorithm 2 Pseudocode of Partial-layer Compression
0:  Original LLM: $M$, number of model layers: $N$, overall compression ratio: $R_{o}$, intermediate rank: $r_{1}$, step: $s$
0:  Number of layers to compress: $k$, layer compression ratio: $R_{l}$
2:  $\mathrm{Set}_{\boldsymbol{W}}\leftarrow M$triangleright Extract the set of weight matrices in $M$
3:  $\mathrm{Set}_{k}\leftarrow\{\,k\mid k=s,2s,\dots,N-s\}$triangleright Obtain candidate compressed layer numbers
4:  $Lowest\_Err\leftarrow+\infty$
5:  for each $k$ in $\mathrm{Set}_{k}$ do
6:     $R_{l}\leftarrow(N\cdot R_{o})/k$triangleright Calculate corresponding layer compression ratio
7:     if $R_{l}\geq 0$ then
8:        $M_{TMP}\leftarrow$ Deep_Copy$(M)$
9:        for $i\in[N-k+1,N]$ do
10:           $W_{i}\leftarrow$ $\mathrm{Set}_{\boldsymbol{W}}(i)$, $S_{i}\leftarrow$ Whitening$(W_{i},CD)$triangleright Initialize weight matrices to compress
11:           $\mathrm{Set}_{\boldsymbol{W^{\prime}}}(i)\leftarrow$ Residual Compensation($W_{i}S_{i}$, $k$, $R_{l}$, $r_{1}$)
12:           Replace weights of layer $i$ in $M_{TMP}$ with $\mathrm{Set}_{\boldsymbol{W^{\prime}}}(i)$
13:        end for
14:        $Err\leftarrow$ Cal_Error($M,M_{TMP},CD,Layer(N)$) $\triangleright$ Calculate the last layer output error
15:        if $Err<Lowest\_Err$ then
16:           $Lowest\_Err\leftarrow Err$triangleright Search for the lowest layer-wise error
17:           $k^{\star}\leftarrow k$
18:           $R_{l}^{\star}\leftarrow R_{l}$
19:        end if
20:     end if
21:  end for
22:  return  $k^{\star},R_{l}^{\star}$
Algorithm 3 Pseudocode of Residual Compensation
0:  Original LLM: $M$, weight matrix: $\boldsymbol{W}\in\mathbb{R}^{m\times n}$, number of model layers: $N$, overall compression ratio: $R_{l}$, step: $s$, intermediate rank: $r_{1}$
0:  Compressed weight matrix set $\mathrm{Set}_{\boldsymbol{W^{\prime}}}(i)$ of layer $i$
2:  $\mathrm{Set}_{\boldsymbol{W}}\leftarrow M$triangleright Extract the sets of weight matrices in $M$ to compress
3:  $\mathrm{Set}_{\boldsymbol{S}}\leftarrow$ Whitening($\mathrm{Set}_{\boldsymbol{W}}$, $CD$), $\mathrm{Set}_{\boldsymbol{W^{\prime}}}\leftarrow\emptyset$triangleright Initialize sets of weight matrices
4:  $k^{\star}$, $R_{l}^{\star}\leftarrow$ Partial-layer Compression($M$, $N$, $R_{o}$, $s$)
5:  $r=m\cdot n\cdot(1-R_{l}^{\star})/(m+n)$, $r_{2}=r-r_{1}$
6:  for $i\in[N-k^{\star}+1,N]$ do
7:     $\boldsymbol{W}_{i}\leftarrow\mathrm{Set}_{\boldsymbol{W}}(i)$, $\boldsymbol{S}_{i}\leftarrow\mathrm{Set}_{\boldsymbol{S}}(i)$triangleright Extract whitening matrices of current weight matrices
8:     $\boldsymbol{U}_{i,r_{1}}$, $\boldsymbol{\Sigma}_{i,r_{1}}$, $\boldsymbol{V}_{i,r_{1}}^{T}$leftarrow $\boldsymbol{W}_{i,r_{1}}$leftarrow SVD_Trunc($\boldsymbol{W}_{i}\boldsymbol{S}_{i}$) $\triangleright$ Apply the first SVD truncation
9:     $\boldsymbol{R}_{i}$leftarrow Cal_Res($\boldsymbol{W}_{i}$, $\boldsymbol{W}_{i,r_{1}}$) $\triangleright$ Calculate residual matrices
10:     $\boldsymbol{U}_{i,r_{2}}$, $\boldsymbol{\Sigma}_{i,r_{2}}$, $\boldsymbol{V}_{i,r_{2}}^{T}$leftarrow $\boldsymbol{R}_{i,r_{2}}$leftarrow SVD_Trunc($\boldsymbol{R}_{i}$) $\triangleright$ Apply the second SVD truncation
11:     $\hat{\boldsymbol{U}}_{i,r_{1}}$leftarrow Mul($\boldsymbol{U}_{i,r_{1}}$, $\sqrt{\boldsymbol{\Sigma}_{i,r_{1}}}$), $\hat{\boldsymbol{V}}_{i,r_{1}}$leftarrow Mul($\sqrt{\boldsymbol{\Sigma}_{i,r_{1}}}$, $\boldsymbol{V}_{i,r_{1}}^{T}$) $\triangleright$ Absorb singular values
12:     $\hat{\boldsymbol{U}}_{i,r_{2}}$leftarrow Mul($\boldsymbol{U}_{i,r_{2}}$, $\sqrt{\boldsymbol{\Sigma}_{i,r_{2}}}$), $\hat{\boldsymbol{V}}_{i,r_{2}}$leftarrow Mul($\sqrt{\boldsymbol{\Sigma}_{i,r_{2}}}$, $\boldsymbol{V}_{i,r_{2}}^{T}$) $\triangleright$ Absorb singular values
13:     $\hat{\boldsymbol{U}}_{i,r}$leftarrow Comb($\hat{\boldsymbol{U}}_{i,r_{1}}$, $\hat{\boldsymbol{U}}_{i,r_{2}}$), $\hat{\boldsymbol{V}}_{i,r}$leftarrow Comb($\hat{\boldsymbol{V}}_{i,r_{1}}$, $\hat{\boldsymbol{V}}_{i,r_{2}}$) $\triangleright$ Combine weight matrices $\mathrm{Set}_{\boldsymbol{W^{\prime}}}$leftarrow ($\hat{\boldsymbol{U}}_{i,r}$, $\hat{\boldsymbol{V}}_{i,r}$) $\cup$ $\mathrm{Set}_{\boldsymbol{W^{\prime}}}$triangleright Append decomposed weight matrices
14:  end for
15:  return  $\mathrm{Set}_{\boldsymbol{W^{\prime}}}$
A.2 Overall Performance of LLaMA-2-7B
We evaluate the perplexity and zero-shot accuracy of LLaMA-2-7B [39] to assess the effectiveness of ResSVD, with results reported in Table 7. On the three language modeling datasets, ResSVD outperforms all baselines across the evaluated compression ratios, except at the 60% compression ratio on WikiText-2 [29]. In particular, compared to the existing best baseline, the relative improvement ranges from 3% to 75%. For compression ratios of 40%, 50%, and 60%, we also include results from the recent method AdaSVD [22], using values reported in their paper due to the lack of an open-source codebase. ResSVD continues to outperform AdaSVD across these settings. Moreover, on zero-shot accuracy across seven common sense reasoning datasets, ResSVD achieves the best performance on most benchmarks, with an average improvement of up to 16%. These results further demonstrate the robustness and overall effectiveness of ResSVD.
Table 7: Overall performance of LLaMA-2-7B [39] compressed by ResSVD and baselines under 20% to 60% compression ratios ("Ratio"), including performance on three language modeling datasets (measured by perplexity ($\downarrow$)) and zero-shot performance on seven common sense reasoning datasets (measured by individual and average accuracy ($\downarrow$)). The best results are marked in bold. Blue arrows within parentheses highlight the relative improvement over the existing second-best baseline. NaN denotes evaluation failure due to numerical instability.
[…]
- Original 5.47 26.84 7.28 0.31 0.69 0.67 0.56 0.40 0.78 0.27 0.53
20% SVD 18208.79 59320.78 27131.56 0.14 0.26 0.49 0.25 0.22 0.52 0.20 0.30
ASVD [44] 9.56 120.74 12.85 0.25 0.59 0.61 0.46 0.32 0.72 0.24 0.45
SVD-LLM [41] 8.37 139.68 20.18 0.23 0.50 0.59 0.39 0.26 0.65 0.23 0.41
ResSVD 7.63 ($\downarrow$9%) 45.37 ($\downarrow$62%) 14.73 ($\downarrow$8%) 0.28 0.61 0.65 0.50 0.35 0.73 0.26 0.48 ($\uparrow$7%)
30% SVD 30373.39 48930.94 36905.54 0.12 0.25 0.49 0.25 0.22 0.51 0.21 0.29
ASVD [44] 984.03 NaN NaN 0.15 0.27 0.51 0.26 0.22 0.53 0.20 0.31
SVD-LLM [41] 10.66 292.90 34.96 0.21 0.42 0.55 0.34 0.22 0.60 0.23 0.37
ResSVD 10.32 ($\downarrow$3%) 73.04 ($\downarrow$75%) 23.28 ($\downarrow$33%) 0.23 0.51 0.62 0.42 0.29 0.68 0.24 0.43 ($\uparrow$ 16%)
40% SVD 39524.00 68829.98 56518.95 0.13 0.26 0.50 0.25 0.21 0.52 0.18 0.29
ASVD [44] NaN NaN NaN 0.15 0.25 0.50 0.26 0.22 0.52 0.18 0.30
SVD-LLM [41] 16.11 717.34 61.96 0.16 0.35 0.55 0.30 0.20 0.57 0.23 0.34
AdaSVD [22] 14.76 304.62 56.98 0.19 0.41 0.58 0.32 0.23 0.58 - -
ResSVD 14.17 ($\downarrow$4%) 136.32 ($\downarrow$55%) 43.19 ($\downarrow$24%) 0.20 0.43 0.57 0.35 0.24 0.63 0.23 0.38 ($\uparrow$12%)
50% SVD 53405.48 39023.05 58547.82 0.15 0.25 0.48 0.25 0.22 0.53 0.18 0.29
ASVD [44] NaN NaN NaN 0.13 0.26 0.50 0.25 0.23 0.50 0.20 0.30
SVD-LLM [41] 27.19 1775.52 129.71 0.14 0.30 0.50 0.28 0.20 0.54 0.23 0.31
AdaSVD [22] 25.58 593.14 113.84 0.15 0.34 0.54 0.29 0.20 0.56 - -
ResSVD 24.26 ($\downarrow$5%) 286.24 ($\downarrow$52%) 100.34 ($\downarrow$12%) 0.14 0.35 0.55 0.31 0.22 0.59 0.22 0.34 ($\uparrow$10%)
60% SVD 65240.23 79002.21 70659.74 0.14 0.25 0.50 0.25 0.23 0.52 0.19 0.30
ASVD [44] NaN 19581.17 NaN 0.15 0.25 0.50 0.25 0.23 0.52 0.12 0.29
SVD-LLM [41] 54.19 3442.74 263.02 0.14 0.26 0.50 0.27 0.20 0.53 0.21 0.30
AdaSVD [22] 60.08 2137.28 294.26 0.12 0.27 0.50 0.27 0.20 0.53 - -
ResSVD 58.88 (-) 991.48 ($\downarrow$54%) 255.70 ($\downarrow$3%) 0.13 0.29 0.52 0.28 0.21 0.55 0.22 0.31 ($\uparrow$3%)
A.3 More Ablation Results
The ablation results for LLaMA-7B [38] under 40%, 50%, and 60% compression ratios are presented in Table 8. A similar trend to that observed at 20% and 30% compression holds here: incorporating both REC and PLC leads to a substantial reduction in perplexity across all settings. Moreover, ResSVD continues to outperform all existing baselines, further validating the effectiveness of these two components under higher compression ratios.
Table 8: Ablation results of Residual Compensation (REC) and Partial-layer Compression (PLC) on LLaMA-7B [38] under compression ratios 40%, 50% and 60%.
(a) REC
[…]
40% SVD 52326.99 59859.41 47643.04
ASVD [44] 5262.11 8806.33 6522.61
SVD-LLM [41] 13.83 57.07 48.47
ResSVD 13.32 56.25 36.61
ResSVD 12.92 46.93 30.51
50% SVD 130388.72 86721.38 79853.46
ASVD [44] 62726.86 117959.06 77773.84
SVD-LLM [41] 24.05 150.58 141.87
ResSVD 23.06 134.91 87.31
ResSVD 22.41 128.80 83.80
60% SVD 52326.99 59859.41 47643.04
ASVD [44] 16221.43 20119.36 16561.39
SVD-LLM [41] 53.20 378.19 310.17
ResSVD 51.03 357.94 249.36
ResSVD 48.67 323.74 260.06
(b) PLC
[…]
40% SVD 52326.99 59859.41 47643.04
ASVD [44] 5262.11 8806.33 6522.61
SVD-LLM [41] 13.83 57.07 48.47
ResSVD 14.33 50.96 46.40
ResSVD 12.92 46.93 30.51
50% SVD 130388.72 86721.38 79853.46
ASVD [44] 62726.86 117959.06 77773.84
SVD-LLM [41] 24.05 150.58 141.87
ResSVD 25.92 147.24 113.76
ResSVD 22.41 128.80 83.80
60% SVD 52326.99 59859.41 47643.04
ASVD [44] 16221.43 20119.36 16561.39
SVD-LLM [41] 53.20 378.19 310.17
ResSVD 60.12 364.38 304.21
ResSVD 48.67 323.74 260.06
A.4 Demonstration of Generated Contents
Table 9 shows the generation contents of models compressed by ResSVD compared to the original model under zero-shot conditions. The results indicate that across various input questions, compressed models consistently produce fluent, coherent, and highly relevant responses. Even at 40%compression ratio, the compressed model can also deliver accurate and relevant answers to common sense questions. These results highlight the effectiveness of ResSVD in achieving substantial compression without compromising performance.
Table 9: Generated contents of LLaMA-7B [38] compressed by ResSVD.
Question Comp. Ratio Answer
What is the responsibility of the AI assistant? - The AI assistant is responsible for understanding the context and taking appropriate action. The AI assistant is responsible for understanding the intent, context, environment and taking appropriate actions. The AI assistant is responsible for interpreting.
20% The AI assistant should be responsible for providing an answer to questions. This means that when asked a question, you should do your best to answer it and not give any information other than what is requested. You should also consider the tone and intent.
30% The AI Assistant’s job is to provide assistance to those who wish for assistance in their lives. It has no duty or obligation to assist those who do not desire assistance, nor does it have any duty to prevent.
40% The AI assistant does not have a specific role. The AI assistant will provide instructions to follow a path, and will give directions as an instruction to follow the path.
What is the universe? Explain this concept. - The Universe is everything that exists. It includes all space, time, matter and energy. It also includes everything which has happened or will happen. The universe began about 13.7 billion years ago (the Big Bang) and will.
20% The universe is the whole of space and time. It consists of all things in space and all things that happen, or have happened. There are many theories on how it began; there were theories that it had no beginning at all.
30% The universe consists of everything in it, including matter and objects, space, the cosmos or outer space, and the galaxy system, as well as other galaxies that make up the Universe. The universe has its own life cycle.
40% In the universe, life is very diverse and complex. The universe is made up of many different objects. These objects are divided into groups. Each group consists of different objects that have specific properties. This group contains the universe, which is large.
Why do cars need fuel to run? - Fuel is the lifeblood of a car. Without it, there’s nothing to power the pistons and no way to drive. A car needs a constant supply of fuel.
20% The fuel in the tank is burned to produce heat, which produces steam that drives a turbine, which turns a crankshaft. The crankshaft turns the wheels of the car, propelling it forward.
30% The fuel is the fuel in a car and it burns and goes up and up. It’s important because it burns the gasoline, which makes it go faster. It’s important because you can get energy out.
40% The fuel gives energy for the cars. The gasoline is used by a motor, which produces heat and air, heating up the engine. This allows the engines to travel faster, allowing them to reach higher speed.
Why do people feel hungry after exercise? - The answer is because the body burns more calories when you are exercising. You may not realize that the body uses up energy to digest food and absorb nutrients from it.
20% Because the body is more efficient at converting food into energy when inactive, so when active it needs to convert food into energy faster. This is why people often get hungry after exercise.
30% The reason people feel hungry after an activity is because they burn energy and energy is metabolised in the body.
40% The reason the body is hungry after a long athletic activity was discovered by scientists in 1976.
A.5 Computation Complexity
ResSVD decomposes the original weight matrix $\boldsymbol{W}\in\mathbb{R}^{m\times n}$ into two low-rank matrices: $\hat{\boldsymbol{W}}_{u}\in\mathbb{R}^{m\times r}$ and $\hat{\boldsymbol{W}}_{v}\in\mathbb{R}^{r\times n}$. The layer compression ratio $R_{l}$ is computed as $R_{l}=1-\frac{(m+n)r}{mn}$. Under a fixed overall compression ratio $R_{o}$, the relationship between $R_{l}$ and $R_{o}$ is given by $R_{l}=\frac{NR_{o}}{k}$, where $N$ is the total number of layers and $k$ denotes the number of last layers to be compressed.
Given an input activation $\boldsymbol{X}\in\mathbb{R}^{n\times m}$, the original output is computed as $\boldsymbol{Y}=\boldsymbol{W}\boldsymbol{X}$. In the compressed model layer, an intermediate state is first computed as $\boldsymbol{I}=\hat{\boldsymbol{W}}_{v}\boldsymbol{X}$, followed by $\boldsymbol{Y}=\hat{\boldsymbol{W}}_{u}\boldsymbol{I}$. The computational complexity of the original model is $N\cdot\mathcal{O}(m^{2}n)$. For the compressed model, the first $(N-k)$ layers remain uncompressed and retain a complexity of $(N-k)\cdot\mathcal{O}(m^{2}n)$, while the compressed $k$ layers incur a cost of $k\cdot\mathcal{O}(m^{2}r+rnm)$. Thus, the total complexity becomes:
$(N-k)\cdot\mathcal{O}(m^{2}n)+k\cdot\mathcal{O}(m^{2}r+rnm).$ (13)
And the rank $r$ is given by:
$r=\frac{mn(1-R_{l})}{m+n}=\frac{mn(k-NR_{o})}{k(m+n)}.$ (14)
Substituting the expression for $r$ into Equation 13, we obtain the simplified total complexity:
$\boxed{N(1-R_{o})\cdot\mathcal{O}(m^{2}n)}$ (15)
Compared to the original computation complexity $N\cdot\mathcal{O}(m^{2}n)$. This indicates that the total computation cost is reduced proportionally to the overall compression ratio. For example, with $R_{o}=40$%, the compressed model requires only 60% of the original computational cost.
A.6 Limitations
Our strategy is to compress the last few layers of LLMs, which consistently leads to performance improvements. However, for simplicity, we currently apply a uniform compression ratio across all weight matrices within each transformer block. Given that MLP and attention layers may exhibit different sensitivities to compression, assigning module-specific compression ratios in a principled way remains a promising direction for future work.
A.7 Broader Impacts.
By enabling the efficient deployment of LLMs on limited hardware, our method has the potential to broaden access to LLMs. However, we also acknowledge that increased accessibility may lower the barrier to misuse, and thus encourage responsible application in practice.
A.8 Performance on VLMs
We also apply ResSVD to the language part of vision-language models (VLMs) such as LLaVA [26], and evaluate its performance on image captioning (Section A.8.1) and visual question answering (Section A.8.2) tasks. All images used in these evaluations are selected from the ImageNet [8] dataset.
A.8.1 Performance of Image captioning
As shown in Figure 6, LLaVA-1.5-7B compressed with ResSVD under 20% compression ratio is still able to produce image captions that are faithful to the visual content. This indicates that the model retains its ability to interpret and describe images accurately, despite parameter reduction.
Figure 6: Image captioning results of LLaVA-1.5-7B [26] compressed by ResSVD under 20% compression ratio. Captions that accurately describe the image content are highlighted in orange.
A.8.2 Performance of Visual Question Answering
Figure 7 presents the visual question answering results of LLaVA-1.5-7B compressed by ResSVD. The model retains strong performance across these question types, indicating that its multi-modal reasoning capability remains intact despite the compression.
Figure 7: Visual question answering outputs generated by LLaVA-1.5-7B [26] compressed using ResSVD under 20% compression ratio. Questions (Q) and model answers (A) are provided, correct answers are highlighted in orange to emphasize answer quality retention.